I need help solving. Thank you
x: 4x + 3a = 5(2b - x)

Answer by stanbon(75887)

4x + 3a = 5(2b - x)
-------
4x + 3a = 10b - 5x
-------
9x = 10b-3a
------
x = (10b-3a)/9
===================
Cheers,
Stan H.
-------------------